OpenStudy (goformit100):

Solve for x. 4x − (2x + 1) = 3 + 2(x − 2)

OpenStudy (goformit100):

@DarlingDare

OpenStudy (goformit100):

@uri

OpenStudy (anonymous):

distribute out everything first.

OpenStudy (anonymous):

This includes the - and the 2 outside of the parentheses....what do you get then?

OpenStudy (uri):

Lol.

OpenStudy (rsadhvika):

No solutions

OpenStudy (anonymous):

Hi

OpenStudy (jiteshmeghwal9):

4x-2x-1=3+2x-4 2x-1=2x-1 no solution :)

OpenStudy (anonymous):

 so.. no solution! :)